-4/x+1=5/x-3+1/x+1
Answer by stanbon(75887)

-4/(x+1) = 5/(x-3) + 1/(x+1)
--------
Multiply thru by (x+1)(x-3) to get:
-4(x-3) = 5(x+1) + (x-3)
-------
-4x+12 = 5x+5+x-3
--------
-4x + 12 = 6x + 2
10x = 10
x = 1
===================
Cheers,
Stan H.
===================